# On the Non-Orientable $3$- and $4$-Genera of a Knot: Connections and Comparisons

### v1

@@ Abstract.
We define a new quantity, the Euler-normalized non-orientable genus, to connect a variety of ideas in the theory of non-orientable surfaces bounded by knots. This quantity is used to reframe non-orientable slice-torus bounds on the non-orientable $4$-genus, to bound below the Turaev genus as a measure of distance to an alternating knot, and to understand gaps between the $3$- and $4$-dimensional non-orientable genera of pretzel knots. Further, we make connections to essential surfaces in knot complements and the Slope Conjecture.
@@ 1. Introduction
This paper seeks to tie together several threads in the theory of non-orientable surfaces bounded by knots using a new quantity called the Euler-normalized non-orientable genus of a knot. This quantity has both $3$- and $4$-dimensional versions. We investigate this quantity in several settings: first, we reframe the meaning of non-orientable slice torus bounds on the non-orientable $4$-genus. Second, we show that the quantity vanishes on alternating knots and, in fact, yields a lower bound on the Turaev genus. Finally, we use it to understand gaps between the $3$- and $4$-dimensional non-orientable genus for pretzel knots and connected sums. Along the way, we will suggest connections to essential surfaces, the colored Jones polynomial, and the Slope Conjecture.
1.1. Non-Orientable Genera of a Knot
To define the Euler-normalized non-orientable genera of a knot $K\subset S^{3}$, we begin by setting notation. A compact smooth embedded surface $F\subset S^{3}$ with $\partial F=K$ is a spanning surface of $K$. Let $\mathcal{F}_{3}(K)$ be the set of all spanning surfaces of $K$. Define the non-orientable $3$-genus, or crosscap number, of $K$ to be
@@ 1.1. Non-Orientable Genera of a Knot
Similarly, a compact smooth properly embedded surface $F\subset B^{4}$ with $\partial F=K$ is a filling of $K$. Let $\mathcal{F}_{4}(K)$ be the set of all fillings of $K$. The (smooth) non-orientable $4$-genus of a knot $K$ is
@@
We are now ready to define our central object of study, the Euler-normalized non-orientable genus. For a spanning surface or filling $F$ of $K$, define its Euler-normalized first Betti numbers to be
@@
where $\sigma(K)$ is the signature of $K$.
A key feature of $\Gamma^{+}(F)$ (resp. $\Gamma^{-}(F)$) is that it is invariant under the addition of a ++-twisted band (resp. a --twisted band), as in Figure 1. Such an addition raises the first Betti number by $1$ and changes the normal Euler number by $\pm 2$, while leaving the isotopy class of the boundary unchanged. We denote by $\operatorname{tw}_{\pm}(F)$ the addition of a positive or negative twisted band to $F$. This feature of $\Gamma^{\pm}(F)$ is the origin of the moniker “Euler-normalized”.
@@
The $\pm$-Euler-normalized non-orientable $n$-genus of a knot $K$, for $n=3,4$, is defined to be
$\hat{\gamma}^{\pm}_{n}(K)=\min\{\Gamma^{\pm}(F)\;:\;F\in\mathcal{F}_{n}(K)\}.$
We also define the average Euler-normalized $n$-genus of $K$ to be
$\overline{\gamma}_{n}(K)=\frac{1}{2}\left(\hat{\gamma}_{n}^{+}(K)+\hat{\gamma}_{n}^{-}(K)\right).$
We will argue below that all of these quantities are fundamentally different from the (ordinary) non-orientable genera.
@@ 1.2. Inspiration from Geography Questions
The definition of the Euler-normalized non-orientable genus arose from explorations into Allen’s geography problem for the non-orientable $4$-genus [3]. For a knot $K$, Allen asked for a description of the set $R_{4}(K)$ of possible pairs $(e,b)$ realized by the normal Euler number and first Betti number, respectively, of a filling of $K$. We may define $R_{3}(K)$ similarly; note that $R_{3}(K)\subset R_{4}(K)$. As depicted in Figure 2, the set $R_{*}(K)$ is constrained by the Gordon-Litherland signature bound, with points of $R(K)$ lying in the wedge
@@
In the geography diagram, we may read off $\gamma_{n}(K)$ as the lowest height of a point in $R_{n}(K)$, whereas $\hat{\gamma}^{+}_{n}(K)$ (resp. $\hat{\gamma}^{-}_{n}(K)$) is the minimum vertical distance from $R_{n}(K)$ to the upward-sloping (resp. downward-sloping) boundary of $W_{\sigma}$. See Sections 2.1, 3.1, and 3.3, below, for a more detailed discussion.
@@
Figure 2. The geography $R_{*}(K)$ is constrained by the Gordon-Litherland signature bound. The Euler-normalized first Betti number measures the vertical distance from the $(e,b)$ coordinates of a surface to the signature bound.
The geography perspective leads to a reframing of many non-orientable slice-torus bounds. For example, we may reinterpret the Heegaard-Floer $\upsilon$ bound on the non-orientable $4$-genus of [38] as in the following theorem, which we shall prove in Section 3.4:
Theorem 1.2.
@@
$\max\left\{\sigma(K)-2\upsilon(K),0\right\}$ $\leq\hat{\gamma}^{+}_{4}(K),$
$\max\left\{2\upsilon(K)-\sigma(K),0\right\}$ $\leq\hat{\gamma}^{-}_{4}(K).$
1.3. Alternating Knots and the Turaev Genus
Setting aside connections to $4$-dimensional bounds for the moment, we next shift to dimension $3$ to completely characterizes when the positive and negative Euler-normalized non-orientable $3$-genera both vanish.
Theorem 1.3.
A knot $K$ is alternating if and only if $\hat{\gamma}^{+}_{3}(K)=0=\hat{\gamma}^{-}_{3}(K)$.
This theorem shows that the Euler-normalized non-orientable $3$-genus can be thought of as a measurement of distance of a knot to the set of alternating knots. We make a connection between the stable non-orientable $3$-genus and another such measurement, the Turaev genus. The Turaev genus $g_{T}$ was developed by Turaev to study the Jones polynomial and the Tait conjectures, and was formalized in [12]; see Section 4.2 for the relevant definitions. As in Theorem 1.3, the Turaev genus of $K$ vanishes if and only if $K$ is alternating. We obtain the following bound.
Theorem 1.4.
For any knot $K$, $\overline{\gamma}_{3}(K)\leq g_{T}(K)$.
As a corollary, we obtain a new family of torus knots with arbitrarily high Turaev genus.
Corollary 1.5.
For $k\geq 1$, $g_{T}(T(2k,2k-1)\geq k-1$.
In fact, since the Turaev genus is realized on an adequate diagram [31], we make the following conjecture, which holds for knots of Turaev genus at most $1$.
Conjecture 1.6.
For an adequate knot $K$, $g_{T}(K)=\overline{\gamma}_{3}(K)$.
The proof of Theorem 1.3 also suggests a connection to the Jones polynomial and the Slope Conjecture. The proof of the forward direction of Theorem 1.3 is reminiscent of Curtis and Taylor’s proof that the largest (resp. smallest) boundary slope of a basic state surface of an alternating knot $K$ is equal to the signature plus twice the maximum (resp. minimum) degree of the Jones polynomial [10]. The surfaces that realize $\hat{\gamma}^{\pm}_{3}(K)$ for an alternating knot $K$ are precisely those that realized Curtis and Taylor’s Jones polynomial bound. This points to a potential connection between $\hat{\gamma}^{\pm}_{3}$ and the Slope Conjecture [16], at least for alternating, and perhaps even adequate, knots.
@@ 1.4. Genus Gaps
Building upon the notion of a slice knot, it is interesting to compare the $3$- and $4$-genera of a knot. We first address the classical question of how large the gap between $\gamma_{3}(K)$ and $\gamma_{4}(K)$ can be. Jabuka and Van Cott proved that $\gamma_{3}(K)-\gamma_{4}(K)$ may be arbitrarily large for families of torus knots [29]. We show that the gap is also arbitrarily large for a family of pretzel knots.
@@
The gap between the Euler-normalized genera $\hat{\gamma}^{\pm}_{3}(P)$ and $\hat{\gamma}^{\pm}_{4}(P)$, on the other hand, is at most $2$, though it is unclear if it is exactly $2$. To find a non-trivial gap between Euler-normalized genera, we instead examine pretzel knots of the form $P_{n}=P(-3,3,n)$ for odd $n\geq 3$. Knots of this form are slice [20]. Hence, we obtain $\hat{\gamma}^{\pm}_{4}(P_{n})=0$. In contrast, the negative Euler-normalized non-orientable $3$-genus of $P_{n}$ is non-vanishing.
@@
For all odd $n\geq 3$, we have $\hat{\gamma}_{3}^{-}(P(-3,3,n))=2$, while $\hat{\gamma}_{3}^{+}(P(-3,3,n))=0$.
@@
We end by showing that the gap between the Euler-normalized non-orientable $3$- and $4$-genera may be arbitrarily large. Specifically, we show that the Euler-normalized non-orientable $3$-genus is additive under connected sum (see Lemma 7.1) and use that fact to prove:
@@
$\hat{\gamma}^{-}_{3}(K_{n})-\hat{\gamma}^{-}_{4}(K_{n})=2n.$
Note that the same result for $\hat{\gamma}^{+}$ holds by taking mirror images.
@@ 1.5. Plan of the Paper
The remainder of the paper is organized as follows: we review non-orientable spanning surfaces and their geography in Section 2. A parallel discussion of non-orientable fillings and their geography appears in Section 3, culminating in a proof of Theorem 1.2. In Section 4, we explore alternating knots (including a proof of Theorem 1.3) and the Turaev genus (including a proof of Theorem 1.4). Next, we examine situations in which $\gamma_{3}(K)>\gamma_{4}(K)$ (Section 5, which includes a proof of Proposition 1.7) and where $\hat{\gamma}^{\pm}_{3}(K)>\hat{\gamma}^{\pm}_{4}(K)$ (Section 6, which includes a proof of Proposition 1.8). We end in Section 7 with a discussion of the behavior of $\hat{\gamma}^{\pm}_{3}$ under connected sum, which yields a proof of Theorem 1.9.
@@ Acknowledgements
We thank David Futer for a series of enlightening discussions about the context and potential implications of this work. We further thank the participants of the Philadelphia Area Contact / Topology (PACT) Seminar for their feedback on a presentation of work in progress.
@@ 2. Spanning Surfaces in $3$ Dimensions
The goal of this section is to describe the framework necessary to investigate the non-orientable spanning surfaces of a knot. We start by reviewing the boundary slope of a spanning surface and by adapting Allen’s geography problem to the $3$-dimensional setting; this material appears in Section 2.1. In Sections 2.2 and 2.3, we present two well-known constructions of spanning surfaces. We finish with a discussion of essential surfaces, their boundary slopes, and their relationship to the $3$-dimensional geography of a knot in Section 2.4.
@@ 2.1. Boundary Slope and 3-Dimensional Geography
A possibly non-orientable spanning surface $F$ of a knot $K\subset S^{3}$ has two natural homological invariants: its first Betti number $b_{1}(F)$ and its boundary slope $s(F)$. To define the boundary slope, we push $K$ a small amount along $F$ to obtain a knot $K^{\prime}$; the boundary slope is then given by
@@
We typically depict the geography of a knot on a $2\mathbb{Z}\times\mathbb{N}$ grid. For example, Figure 4 shows the $3$-dimensional geography of the trefoil knot. We will verify the correctness of the figure in the next section.
@@
2.2. Construction: State Surfaces
The state surface construction generalizes both the Seifert algorithm and checkerboard surfaces; see [2] or [32], for example. To set notation, we recall the definition of a state of a knot diagram. First, we define a resolution for a crossing of an unoriented link diagram $D$ to be a choice of splitting the crossing into two separate segments in one of two possible ways: an $A$-resolution or a $B$-resolution, as shown in Figure 5. A state $s$ of a knot diagram $D$ is the configuration of disjoint, possibly nested, simple closed curves (called state circles) in the plane that is produced by applying either an $A$- or $B$-resolution at every crossing of $D$. Let $A(s)$ (resp. $B(s)$) be the number of $A$ resolutions (resp. $B$ resolutions) in the state $s$.
[…]
Figure 5. A crossing of a knot diagram can be split with an $A$-resolution or a $B$-resolution.
To associate a surface to a state $s$, fill each state circle with a disk, thinking of disks of inner nested state circles as lying above those of outer nested state circles in $\mathbb{R}^{3}$. Connect the disks with half-twisted bands where the crossings were. We say that a state surface is basic if no bands connect a state circle to itself. Examples of basic state surfaces for the $5_{2}$ knot are given below in Figure 6.
State surfaces play a particularly important role for the $3$-dimensional geography of alternating knots due to the following result of Adams and Kindred.
Theorem 2.2 (Theorem 5.4 of [2]).
Given a diagram $D$ and a spanning surface $F$ of a knot $K$, there exists a basic state surface $F_{s}$ of the same orientability as $F$ and a (possibly empty) set of twisted bands transforming $F_{s}$ into $F^{\prime}_{s}$ so that $b_{1}(F^{\prime}_{s})=b_{1}(F)$ and $e_{3}(F^{\prime}_{s})=e_{3}(F)$.
This theorem essentially says that state surfaces generate the $3$-dimensional geography of a knot. In particular, we can refine Equation (2.2) for alternating knots as follows.
Corollary 2.3.
Given a diagram $D$ of a knot $K$, let $\mathcal{S}_{n}$ be the set of non-orientable basic state surfaces for $D$, let $\operatorname{tw}_{\pm}(\mathcal{S}_{o})$ be the set of orientable basic state surfaces with one $\pm$-twisted band attached, and finally let $\mathcal{S}=\mathcal{S}_{n}\cup\operatorname{tw}_{+}(\mathcal{S}_{o})\cup\operatorname{tw}_{-}(\mathcal{S}_{o})$. Then
$R_{3}(K)=\bigcup_{F\in\mathcal{S}}W_{F}.$
The normal Euler number of a state surface may be computed combinatorially. We denote by $\operatorname{wr}(D)$ the writhe of a knot diagram $D$.
Lemma 2.4 (Proposition 3.2 of [2]).
If $F_{s}$ is a state surface arising from a state $s$ of a diagram $D$, then [^1] There is a sign error in [2, Proposition 3.2], which is corrected in the statement of Lemma 2.4.
$e_{3}(F_{s})=A(s)-B(s)-\operatorname{wr}(D).$
Example 2.5.
Of the eight state surfaces for the diagram of the trefoil in Figure 3, exactly two are basic. These basic state surfaces arise from the all-$A$ state $s_{A}$, which yields the Seifert surface, and the all-$B$ state $s_{B}$, which yields the Möbius strip in Figure 3. By Lemma 2.4, we compute
$e_{3}(F_{s_{A}})$ $=0$ $e_{3}(F_{s_{B}})$ $=-6$
$b_{1}(F_{s_{A}})$ $=2$ $b_{1}(F_{s_{B}})$ $=1.$
Corollary 2.3 implies that the geography of the trefoil is generated by wedges based at (6,1)(-6,1) (from $F_{s_{B}}$), (2,3)(-2,3), and $(2,3)$ (from adding twists to the Seifert surface $F_{s_{A}}$). This justifies the geography depicted in Figure 4.
Example 2.6.
Of the $32$ state surfaces for the diagram of the $5_{2}$ knot in Figure 6, only three are basic: the all-$A$ state surface $F_{s_{A}}$, the all-$B$ state surface $F_{s_{B}}$, and one more state surface $F_{s}$; see Figure 6.
[…]
Figure 6. The three basic state surfaces for the diagram of the $5_{2}$ knot at top.
As above, we compute that
$e_{3}(F_{s_{A}})$ $=10$ $e_{3}(F_{s})$ $=4$ $e_{3}(F_{s_{B}})$ $=0$
$b_{1}(F_{s_{A}})$ $=3$ $b_{1}(F_{s})$ $=2$ $b_{1}(F_{s_{B}})$ $=2.$
The all-$B$ state surface is orientable, and hence Corollary 2.3 implies that the geography of the $5_{2}$ knot is generated by wedges based at $(10,3)$, $(4,2)$, $(2,3)$, and (2,3)(-2,3). See Figure 7.
[…]
Figure 7. The 33-dimensional geography of the $5_{2}$ knot is generated by the state surfaces for $s_{A}$ at $(10,3)$ and $s$ at $(4,2)$, along with the state surface $s_{B}$ (which is depicted by an open dot at $(0,2)$) with a negative twisted band at (2,3)(-2,3).
2.3. Construction: Pinch Surfaces for Torus Knots
In this section, we describe the construction of a pinch surface of a torus knot $T(p,q)$ in the case that $pq$ is even. This construction is due to Batson [5] for surfaces in $B^{4}$ and adapted to produce surfaces in $S^{3}$ in [29]; we follow the latter source here.
Working abstractly at first, realize $T(p,q)$ as a straight line on a flat torus as in Figure 8(a). Perform a (non-orientable) band move on $T(p,q)$ along a band that connects two adjacent strands, as in Figure 8(b), yielding a new torus knot $T(r,s)$. The parameters $r,s$ satisfy [5, 30]
(2.3) $r$ $=|p-2t|$ $s$ $=|q-2h|$
where $t\equiv-q^{-1}\mod p$ and $h\equiv p^{-1}\mod q$.
[…]
Figure 8. (a) A non-orientable band for $T(p,q)$, (b) performing a band move, and (c) isotoping the resulting knot to obtain $T(r,s)$.
More concretely, place $T(p,q)$ on a standard torus in $\mathbb{R}^{3}$ so that it wraps $p$ times around the longitude and $q$ times around the meridian. Place $T(r,s)$ on a torus contained in the interior of the first. The band move above creates an embedded cobordism between the two knots.
Since the parity of $pq$ is preserved by pinch moves [29, Lemma 2.1], repeated pinching of $T(p,q)$ yields the unknot $T(k,1)$ with $k$ even. Further pinch moves reduce $k$ by $2$, eventually yielding the meridian $T(0,1)$. Capping the meridian by a disk finishes the construction of the $3$-dimensional pinch surface $F_{3}(p,q)$.
As pointed out in [29], the pinch surface realizes the non-orientable $3$-genus of $T(p,q)$, which was computed by Teregaito [42]. The first Betti number of $F_{3}(p,q)$ is given by the number of pinches. Further, we may compute the normal Euler number of the pinch surface.
Lemma 2.7 (Lemma 2.3 of [39] or Proposition 6 of [42]).
The pinch surface $F_{3}(p,q)$ has normal Euler number $e_{3}(F_{3}(p,q))=-pq$.
Example 2.8.
The pinch surface for the trefoil knot $T(3,2)$ coincides with the Möbius band in Figure 3. The result of Lemma 2.7, namely that $e_{3}(F_{3}(3,2))=-6$, agrees with the computations in Sections 2.1 and $2.$
Remark 2.9.
There is an extension of the pinch surface construction to the case where $pq$ is odd that involves splitting $T(p,q)$ into a pair of even torus knots on the boundaries of complementary solid tori, and then applying the algorithm above to each of the resulting knots; see [29] for details.

### v2

# A Refinement of the Spanning Surface Defect in $3$ and $4$ Dimensions

@@ Abstract.
The spanning surface defect uses spanning surfaces of a knot in the $3$-sphere to measure how far a knot is from being alternating. We refine the spanning surface defect and extend the definition to take into account surfaces in the $4$-ball. We use these extensions to make comparisons between the $3$- and $4$-dimensional settings, to reframe non-orientable slice-torus bounds on the non-orientable $4$-genus, and to prove a connected sum formula.
@@ 1. Introduction
Surfaces that span a knot in the $3$-sphere — that is, compact embedded surfaces whose boundary is a given knot — are important tools in understanding the topology of the complement of a knot. For example, one might seek to understand the minimal genus of an orientable — or non-orientable — spanning surface of a knot, or to enumerate the set of possible slopes that spanning surfaces imprint on the boundary of a tubular neighborhood of a knot, or even to characterize whether a knot has an alternating diagram using particularly nice spanning surfaces [17, 22]. In extending the ideas of Greene’s and Howie’s characterizations of alternating knots to a wider class of knots, Ito [26] defined the spanning surface defect to measure how far a knot is from being alternating. The goal of this paper is to refine this measurement and to extend it to $4$ dimensions.
1.1. Spanning Surfaces, Fillings, and the Defect
To define and refine the spanning surface defect of a knot $K\subset S^{3}$, we begin by setting notation. A compact smooth embedded surface $F\subset S^{3}$ with $\partial F=K$ is a spanning surface of $K$. Let $\mathcal{F}_{3}(K)$ be the set of all spanning surfaces of $K$. Define the non-orientable $3$-genus, or crosscap number, of $K$ to be
@@ 1.1. Spanning Surfaces, Fillings, and the Defect
Similarly, a compact smooth properly embedded surface $F\subset B^{4}$ with $\partial F=K$ is a filling of $K$. Let $\mathcal{F}_{4}(K)$ be the set of all fillings of $K$. The non-orientable $4$-genus of a knot $K$ is
@@
We are now ready to define our central object of study, the spanning surface defect. In fact, we will define several flavors of the defect. First, we normalize the genus of a spanning surface or filling $F$ of $K$ using the normal Euler number of $F$ and the signature of $K$. In particular, for a spanning surface or filling $F$ of $K$, define its Euler-normalized first Betti numbers to be
@@
where $\sigma(K)$ is the signature of $K$. Note that the quantity $\sigma(K)-\frac{1}{2}e(F)$ is the signature of the Gordon-Litherland form of $F$; see Section 3.2, below.
A key feature of $\Gamma^{+}(F)$ (resp. $\Gamma^{-}(F)$) is that it is invariant under the addition of a ++-twisted band (resp. a --twisted band), as in Figure 1. Such a band addition raises the first Betti number by $1$ and changes the normal Euler number by $\pm 2$, while leaving the isotopy class of the boundary unchanged. We denote by $\operatorname{tw}_{\pm}(F)$ the addition of a $\pm$-twisted band to $F$. This feature of $\Gamma^{\pm}(F)$ is the origin of the moniker “Euler-normalized”.
@@
The $\pm$-spanning surface defect in dimension $n$ of a knot $K$, for $n=3,4$, is defined to be
$d^{\pm}_{n}(K)=\min\{\Gamma^{\pm}(F)\;:\;F\in\mathcal{F}_{n}(K)\}.$
It is straightforward to check that $\frac{1}{2}(d_{3}^{+}(K)+d_{3}^{-}(K))$ coincides with Ito’s original definition of the spanning surface defect [26]:
$d_{3}(K)=\frac{1}{2}\min\left\{b_{1}(F)+b_{1}(G)-|\sigma_{gl}(F)-\sigma_{gl}(G)|\;:\;F,G\in\mathcal{F}_{3}(K)\right\},$
where $\sigma_{gl}$ is the signature of the Gordon-Litherland form of a spanning surface. Ito’s motivation for defining $d_{3}$ is encapsulated by the following theorem, which we refine slightly in light of Definition 1.1.
Theorem 1.2 ([17, 22] via [26]).
A knot $K$ is alternating if and only if $d^{+}_{3}(K)=0=d^{-}_{3}(K)$.
Ito further shows that the spanning surface defect is a lower bound on the alternating genus of a knot, and hence on the Turaev genus. The Turaev genus $g_{T}$ was developed by Turaev to study the Jones polynomial and the Tait conjectures, and was formalized in [10]. As in Theorem 1.2, the Turaev genus of $K$ vanishes if and only if $K$ is alternating. We obtain the following bound.
Theorem 1.3 ([26] and [34]).
For any knot $K$, $d_{3}(K)\leq g_{T}(K)$.
In fact, since the Turaev genus is realized on an adequate diagram [30], we make the following conjecture.
Conjecture 1.4.
For an adequate knot $K$, $g_{T}(K)=d_{3}(K)$.
@@ 1.2. Inspiration from Geography Questions
Our exploration of the $\pm$-spanning surface defect arose from Allen’s geography problem for the non-orientable $4$-genus [3]; Howie has also taken this perspective for spanning surfaces in $S^{3}$ [23]. For a knot $K$, Allen asked for a description of the set $R_{4}(K)$ of possible pairs $(e,b)$ realized by the normal Euler number and first Betti number, respectively, of a filling of $K$. We may define $R_{3}(K)$ similarly; note that $R_{3}(K)\subset R_{4}(K)$. As depicted in Figure 2, the set $R_{*}(K)$ is constrained by the Gordon-Litherland signature bound, with points of $R(K)$ lying in the wedge
@@
In the geography diagram, we may read off $\gamma_{n}(K)$ as the lowest height of a point in $R_{n}(K)$, whereas $d^{+}_{n}(K)$ (resp. $d^{-}_{n}(K)$) is the minimum vertical distance from $R_{n}(K)$ to the upward-sloping (resp. downward-sloping) boundary of $W_{\sigma}$.
@@
Figure 2. The geography $R_{*}(K)$ is constrained by the Gordon-Litherland signature bound. The $\pm$-spanning surface defect measures the vertical distance from the $(e,b)$ coordinates of a surface to the signature bound.
The discussion above suggests that it may be profitable to think about the $\pm$-spanning surface defects as a type of non-orientable genus normalized by the signature and normal Euler number. For example, we may reinterpret the Heegaard-Floer $\upsilon$ bound on the non-orientable $4$-genus of [37] as in the following proposition, which we shall prove in Section 3.3:
Proposition 1.5.
@@
$\max\left\{\sigma(K)-2\upsilon(K),0\right\}$ $\leq d^{+}_{4}(K),$
$\max\left\{2\upsilon(K)-\sigma(K),0\right\}$ $\leq d^{-}_{4}(K).$
@@ 1.3. Genus Gaps
As in the orientable case, it is interesting to compare the non-orientable $3$- and $4$-genera of a knot. We first address the question of how large the gap between $\gamma_{3}(K)$ and $\gamma_{4}(K)$ can be. Jabuka and Van Cott proved that the gap $\gamma_{3}(K)-\gamma_{4}(K)$ may be arbitrarily large for families of torus knots [28]. We show that the gap is also arbitrarily large for a family of pretzel knots.
@@
The gap between the spanning surface defects $d^{\pm}_{3}(P)$ and $d^{\pm}_{4}(P)$, on the other hand, is at most $2$, though it is unclear if it is exactly $2$. To find a non-trivial gap between spanning surface defects, we instead examine pretzel knots of the form $P_{n}=P(-3,3,n)$ for odd $n\geq 3$. Knots of this form are slice [16]. Hence, we obtain $d^{\pm}_{4}(P_{n})=0$. In contrast, the negative Euler-normalized non-orientable $3$-genus of $P_{n}$ is non-vanishing.
@@
For all odd $n\geq 3$, we have $d_{3}^{-}(P(-3,3,n))=2$, while $d_{3}^{+}(P(-3,3,n))=0$.
@@
We end by showing that the gap between the Euler-normalized non-orientable $3$- and $4$-genera may be arbitrarily large. Specifically, we show that the Euler-normalized non-orientable $3$-genus is additive under connected sum (see Lemma 7.1); we note that Josh Howie independently proved this proposition [23]. We then use the connected sum formula to prove:
@@
$d^{-}_{3}(K_{n})-d^{-}_{4}(K_{n})=2n.$
The same result for $d^{+}$ holds by taking mirror images.
@@ 1.4. Plan of the Paper
The remainder of the paper is organized as follows: we review non-orientable spanning surfaces and their geography in Section 2. A parallel discussion of non-orientable fillings and their geography appears in Section 3, culminating in a proof of Proposition 1.5. Next, we examine situations in which $\gamma_{3}(K)>\gamma_{4}(K)$ (Section 5, which includes a proof of Proposition 1.6) and where $d^{\pm}_{3}(K)>d^{\pm}_{4}(K)$ (Section 6, which includes a proof of Proposition 1.7). We end in Section 7 with a discussion of the behavior of $d^{\pm}_{3}$ under connected sum, which yields a proof of Theorem 1.8.
@@ Acknowledgements
We thank David Futer and Josh Howie for enlightening discussions about the context and potential implications of this work. We further thank the participants of the Philadelphia Area Contact / Topology (PACT) Seminar for their feedback on a presentation of work in progress.
@@ 2. Spanning Surfaces in $3$ Dimensions
The goal of this section is to describe the framework necessary to investigate the non-orientable spanning surfaces of a knot. We start by reviewing the boundary slope of a spanning surface and by adapting Allen’s geography problem to the $3$-dimensional setting; again, Howie has also taken this perspective. We finish with a discussion of essential surfaces, their boundary slopes, and their relationship to the $3$-dimensional geography of a knot in Section 2.2.
@@ 2.1. Boundary Slope and 3-Dimensional Geography
A possibly non-orientable spanning surface $F$ of a knot $K\subset S^{3}$ has two natural homological invariants: its first Betti number $b_{1}(F)$ and its boundary slope $s(F)$. To define the boundary slope, we push $K$ by a small amount along $F$ to obtain a knot $K^{\prime}$; the boundary slope is then given by
@@
We typically depict the geography of a knot on a $2\mathbb{Z}\times\mathbb{N}$ grid. For example, Figure 4 shows the $3$-dimensional geography of the trefoil knot; we will discuss why the figure is correct at the end of this subsection.
@@
To see that Figure 4 correctly depicts the $3$-dimensional geography of the trefoil knot, we invoke Adams and Kindred’s work on state surfaces of alternating knots [2]. They show that the $3$-dimensional geography of an alternating knot is the union of the wedges of the non-orientable basic state surfaces of a reduced alternating diagram, together with the wedges of the orientable basic state surfaces with twisted bands added. Since the only basic state surfaces for the diagram of the trefoil in Figure 3 are the Seifert surface and the Möbius strip, we obtain the geography in Figure 4.